OpenStudy (anonymous):

how to solve 4x+2=3x-8?

OpenStudy (anonymous):

Put all the x related numbers on one side of the equation and all the rest on the other side of the equality sign. Then do your subtractions additions such that now you only have one x in the whole question. Then you isolate x by division.

OpenStudy (anonymous):

4x + 2 = 3x - 8 -3x -3x -------------- x + 2 = -8 - 2 - 2 -------------- x = -10 What you must do is isolate the variable on one side of the equation; to do that, you must use the opposite operand(+ with -, * with /).

OpenStudy (anonymous):

Thank you. I did what the first person told me and I knew it was wrong. I did this and got the same answer.

OpenStudy (anonymous):

hhmmm... sorry if I didn't make my explanation clearer.